OpenStudy (anonymous):

Solve using the square root method. Round the solution to the nearest tenth. x2 - 20 = 0 A. {-10.0, 10.0} B. {-5.5, 5.5} C. {-3.5, 3.5} D. {-4.5, 4.5} Solve using the square root method. Round the solution to the nearest tenth. x2 - 20 = 0 A. {-10.0, 10.0} B. {-5.5, 5.5} C. {-3.5, 3.5} D. {-4.5, 4.5} @Mathematics

OpenStudy (anonymous):

See the other reply to your question and this problem should become rather easy to solve.

jhonyy9 (jhonyy9):

x2 =4x5 x_1,2=+/- 2sqrt5

OpenStudy (anonymous):

what in the world is the answer to this?

OpenStudy (anonymous):

\[x^2 = 20\] \[x = \pm \sqrt{20}\] \[x = \pm \sqrt{4}\sqrt{5}\] \[ x = \pm 2\sqrt{5}\] \[x_1 = 2\sqrt{5} \approx 4.5\] \[x_2 = -2\sqrt{5} \approx -4.5\] so the right answer is D

OpenStudy (anonymous):

There it is...haha no calculator so I didn't compute 2root5